OpenStudy (anonymous):

The price f(x), in dollars, of product A after x years is represented by the function below: f(x) = 72(1.25)x Part A: Is the price of product A increasing or decreasing and by what percentage per year? Justify your answer. (5 points) Part B: The table below shows the price f(t), in dollars, of product B after t years: t (number of years) 1 2 3 4 f(t) (price in dollars) 65 84.5 109.85 142.81 Which product recorded a greater percentage change in price over the previous year?

OpenStudy (mathmale):

Please do yourself and others a favor by expressing exponentiation properly. f(x) = 72(1.25)x is incorrect; you mean\[f(x) = 72(1.25)^x.\]

OpenStudy (mathmale):

Evaluate 1.25^x at 0, 1, 2, 3. Is this factor increasing or decreasing or remaining the same?

OpenStudy (anonymous):

I copied and pasted the question so idk what you are talking about.. ._.

OpenStudy (anonymous):

and its decreasing

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Copying and pasting does not always put the exponent (x) in the correct place. 75(1.25)^x denotes exponentiation: It's 1.25, raised to the power x, and then the result is multiplied by 75. As before, I ask you to evaluate 1.25^x for x=0, 1, 2, 3. Is this function increasing or decreasing as x increases?

OpenStudy (anonymous):

idk.. @mathmale

OpenStudy (mathmale):

(any number other than zero)^0=1. Therefore, 1.25^0 = ? 1.25^1 = ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

1.25^0 does not equal 0. That 0 is an exponent. Note that 2^0 = 1. 3^0 =1. What is 1.25^0?

OpenStudy (anonymous):

1? @mathmale

OpenStudy (mathmale):

Yes, that's right. What is 1.25^1? 1.25^2? iF YOU'RE

OpenStudy (mathmale):

If you're not sure, review exponentiation, please.

OpenStudy (anonymous):

Ok so whats next? @mathmale

OpenStudy (mathmale):

Have you found 1.25^1, 1.25^2 and 1.25^3 yet?

OpenStudy (mathmale):

As x increases, does 1.25^x increase, decrease or stay the same?

OpenStudy (mathmale):

If you can answer this correctly, you will have essentially answered part A of your question.

OpenStudy (mathmale):

B: Sorry this discussion has dragged on and on. But if you go back and review it, you'll see that there were several occasions on which you didn't follow through. For example, I asked you to evaluate 1.25^1, 1.25^2 and 1.25^3. Hope we can finish up this problem later on.